What is the change in the gravitational potential energy of a Boeing 767 jet as it soars from the runway up to a cruising altitude of 10.2 km? Assume its mass is a constant 2.04×10[sup]5[/sup] kg. answer in joules.
1
Expert's answer
2011-03-09T06:33:36-0500
The change in the gravitational potential energy would be equal to
Δ P = mg Δ H = 2.04×105 [kg]*10 [ms-2]* 10.2 ×103 [m] = 20.808 × 109 J.